OpenStudy (anonymous):

subtract and simplify: (3x/x-5) - (5/5-x)

OpenStudy (precal):

post your equation please

OpenStudy (anonymous):

\[3x/x-5 - 5/5-x\]

OpenStudy (shane_b):

I feel parentheses would help here...

OpenStudy (anonymous):

okay i re-wrote it on the top.

OpenStudy (precal):

yes we need ( ) and to solve it would be equal to something, unless the instructions were to simplify

OpenStudy (anonymous):

x-3 ?

OpenStudy (anonymous):

I am assuming that it's \[\frac{3x}{x - 5} - \frac{5}{5 - x}\] \[\frac{3x}{x - 5} - \frac{5}{5 - x} = \frac{3x}{x - 5} + \frac{5}{x - 5}\]Now just combine the numerators.\[\frac{?}{x - 5}\]

mathslover (mathslover):

\[\huge{\frac{3x}{x-5}-\frac{5}{5-x}}\] Is this the question ? right

OpenStudy (anonymous):

yes

mathslover (mathslover):

ok ! can u write the above as : \[\huge{\frac{3x}{x-5}+\frac{5}{x-5}}\] \[\huge{\frac{3x+5}{x-5}}\] That is all

OpenStudy (anonymous):

okay. thank you! :)

mathslover (mathslover):

best of luck : no problem